Use a suitable identity to get each of the following products.
(7a - 9b) (7a - 9b)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(7a-9b)(7a-9b)= (7a-9b)^2
= 49a^2- 126ab + 81b^2
Using (a-b)^2= a^2+ b^2- 2ab

Thanks